OpenStudy (anonymous):

I NEED HELP SOLVING THIS!! SOMEONE PLEASE HELP!! Which equation can be used to solve the following word problem? Jessica has 9 more nickels than dimes, and the total value of her coins is $1.50. How many nickels does she have? Let n represent the number of nickels. 0.10n + 0.05n + 9 = 1.50 0.10(n – 9) + 0.05n = 1.50 0.10n + 0.05(n + 9) = 1.50 0.10(n + 9) + 0.05n = 1.50

OpenStudy (anonymous):

@jenny171

OpenStudy (anonymous):

@countryboy212

OpenStudy (anonymous):

B

OpenStudy (anonymous):

What algebraic property does this statement show? 3 + (7 + 2) = (3 + 7) + 2 (Points : 5) associative property<< closure property commutative property symmetric property Question 2.2. Solve. 2x – 7 = 6x – 15 (Points : 5) 2<< 3 5 6 Question 3.3. What is the slope of the line that passes through the points (1, 3) and (5, –2)? (Points : 5) Question 4.4. The sum of three consecutive integers is –27. What are the numbers? (Points : 5) {–5, –9, –13} {–10, –9, –8} {–7, –9, –11} {–11, –13, –15} (I need help on this one, I didn’t understand it.) Question 5.5. Which equation can be used to solve the following word problem? Jessica has 9 more nickels than dimes, and the total value of her coins is $1.50. How many nickels does she have? Let n represent the number of nickels. (Points : 5) 0.10n + 0.05n + 9 = 1.50 0.10(n – 9) + 0.05n = 1.50<< 0.10n + 0.05(n + 9) = 1.50 0.10(n + 9) + 0.05n = 1.50 Question 6.6. If A = {2, 4, 6, 8, 10} and B = {1, 2, 3, 4, 5}, find A 09_04_ST_15.gif B. (Points : 5) {1, 2, 3, 4, 5, 6, 8, 10} {1, 3, 5} {2, 4}<< {7, 9} Question 8.8. Select the pair of equations whose graphs are perpendicular. (Points : 5) A: 5x – 8y = 9 12x – 5y = 7 B: 3x + 6y = 8 y = 3x – 8 C: y = 2x – 7 x + 3y = 3 D: 2y = –3x + 5<< 2x – 3y = 4 Question 9.9. Combine like terms to simplify the expression: 7x + 3y – 2 + 6x – 1 + y2 (Points : 5) 2y3 + 13x – 3 3y3 + 13x2 – 3 y2 + 3y + 13x – 3<< y2 + 3y + 13x2 – 3 Question 11.11. Given the linear equation x – y = 5, find the slope of its graph. (Points : 5) –5 –1<< 1 5 Question 13.13. Evaluate the expression when a = –1.4 and b = –2.7. |a| – b (Points : 5) 4.1<< –1.3 –4.1 1.3 Question 16.16. Solve the system of equations. 3x + 3y = –6 –x + y = 0 (Points : 5) x = –1; y = –1<< x = 10; y = –3 x = 22; y = 7 x = 14; y = 5

OpenStudy (anonymous):

@jenny171

OpenStudy (anonymous):

These were the ones lol @jenny171

OpenStudy (anonymous):

3. -5/4

OpenStudy (anonymous):

Is that the only one I got wrong?

OpenStudy (anonymous):

@jenny171

OpenStudy (anonymous):

11: 1

OpenStudy (anonymous):

13: -1.3

OpenStudy (anonymous):

Can you check my answers @Nnesha please?

Nnesha (nnesha):

omg that's a lot :P which one i have to check it ??

OpenStudy (anonymous):

I put the answers in "<<" so you know I put that as my answer

OpenStudy (anonymous):

lol @Nnesha sorry to bother you

Nnesha (nnesha):

replace x and y by their answer choices you got if you get both side equal then ur answer is right simple :)

Nnesha (nnesha):

and u can use calculator ;) to check ur answer

OpenStudy (anonymous):

Which one are you talking about?

Nnesha (nnesha):

wait you tagged me bec i'm human calculator ^^:P:P okay then wait

OpenStudy (anonymous):

lol xD @Nnesha

Nnesha (nnesha):

okay so last one  replace x and y by their values and then solve left side if both side equal then that is ur answer